# EmotionTalk: An Interactive Chinese Multimodal Emotion Dataset With Rich Annotations

### v1

@@ 3.3 Statistics
For the video modality, we separately analyzed the emotion recognition results for videos without and with audio, as shown in Fig 1(c) and 1(d), respectively. In the video modality with audio, neutral emotions constitute the largest proportion, accounting for 6,608 instances (34.3%). Happy and angry emotions also represent significant portions, with 4,451 (23.1%) and 4,190 (21.8%) instances, respectively. The remaining emotional categories include disgusted (1,563 instances, 8.1%), surprised (1,084 instances, 5.6%), sad (964 instances, 5.0%), and fearful (390 instances, 2.0%). Overall, the emotional distribution is notably concentrated in the neutral, happy, and angry categories.
In the video modality, neutral emotions account for 6,853 instances, representing 35.6% of the total. Angry and happy emotions are the next most prevalent, with 3,698 instances (19.2%) and 3,592 instances (18.7%), respectively. The remaining categories include surprised (1,661 instances, 8.6%), disgusted (1,371 instances, 7.1%), sad (1,153 instances, 6.0%), and fearful (922 instances, 4.8%). Overall, the distribution is relatively balanced, with neutral, angry, and happy emotions comprising the majority.
As for the multi-modality, we analyzed the annotation results of continuous emotion labels. The data shows a clear dominance of negative sentiments, with weakly negative and negative emotions making up 78.4% of the total (49.1% for weakly negative and 29.3% for negative). Neutral sentiments account for 11.5%, while positive sentiments (comprising weakly positive and positive) are limited to 10.2%, with weakly positive at 8.2% and positive at 2.0%. This distribution highlights a noticeable bias toward negative emotional expressions in the dataset, which may reflect the inherent characteristics of the video content or the nature of user responses.

### v2

@@ 3.3 Statistics
For the video modality, we separately analyzed the emotion recognition results for videos without and with audio, as shown in Fig 1(c) and 1(d), respectively. In the video modality without audio, neutral emotions constitute the largest proportion, accounting for 6,608 instances (34.3%). Happy and angry emotions also represent significant portions, with 4,451 (23.1%) and 4,190 (21.8%) instances, respectively. The remaining emotional categories include disgusted (1,563 instances, 8.1%), surprised (1,084 instances, 5.6%), sad (964 instances, 5.0%), and fearful (390 instances, 2.0%). Overall, the emotional distribution is notably concentrated in the neutral, happy, and angry categories.
For the multi-modal video modality incorporating audio, neutral emotions constitute the largest category, with 6,853 instances, representing 35.6% of the dataset. Subsequently, angry and happy emotions are the next most frequent, occurring in 3,698 instances (19.2%) and 3,592 instances (18.7%), respectively. The remaining emotional categories are surprised (1,661 instances, 8.6%), disgusted (1,371 instances, 7.1%), sad (1,153 instances, 6.0%), and fearful (922 instances, 4.8%). In summary, the distribution is characterized by a significant proportion of neutral, angry, and happy emotions.
Concerning the continuous emotion labels within the multi-modal data, our analysis demonstrates a marked dominance of negative sentiments. Weakly negative and negative emotions together comprise 78.4% of the total (49.1% and 29.3%, respectively). Neutral sentiments account for 11.5%, while positive sentiments (weakly positive at 8.2% and positive at 2.0%) represent a smaller portion at 10.2%. This distribution underscores a notable bias toward negative emotional expressions in the dataset, which may be attributed to the inherent characteristics of the video content or the nature of user responses.